OpenStudy (anonymous):

I can't find the value of x please help: -3(3x + 15) - (10 + x) = 35

OpenStudy (anonymous):

-3(3x+15)-(10+x)=35 -9x-45-10-x=35 -10x-55=35 -10x=20 x=-20/10 x=-2

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

No prob